# UniGaussian: Driving Scene Reconstruction from Multiple Camera Models via Unified Gaussian Representations

### v1

@@ Abstract
Urban scene reconstruction is crucial for real-world autonomous driving simulators. Although existing methods have achieved photorealistic reconstruction, they mostly focus on pinhole cameras and neglect fisheye cameras. In fact, how to effectively simulate fisheye cameras in driving scene remains an unsolved problem. In this work, we propose UniGaussian, a novel approach that learns a unified 3D Gaussian representation from multiple camera models for urban scene reconstruction in autonomous driving. Our contributions are two-fold. First, we propose a new differentiable rendering method that distorts 3D Gaussians using a series of affine transformations tailored to fisheye camera models. This addresses the compatibility issue of 3D Gaussian splatting with fisheye cameras, which is hindered by light ray distortion caused by lenses or mirrors. Besides, our method maintains real-time rendering while ensuring differentiability. Second, built on the differentiable rendering method, we design a new framework that learns a unified Gaussian representation from multiple camera models. By applying affine transformations to adapt different camera models and regularizing the shared Gaussians with supervision from different modalities, our framework learns a unified 3D Gaussian representation with input data from multiple sources and achieves holistic driving scene understanding. As a result, our approach models multiple sensors (pinhole and fisheye cameras) and modalities (depth, semantic, normal and LiDAR point clouds). Our experiments show that our method achieves superior rendering quality and fast rendering speed for driving scene simulation.
Figure 1: An illustration of the proposed approach. UniGaussian reconstructs 3D driving scenes by learning a unified 3D Gaussian representation from multiple input sources. It achieves holistic driving scene understanding and models multiple sensors (pinhole cameras and fisheye cameras) and modalities (semantic, normal, depth and optional LiDAR point clouds).
@@ 1 Introduction
Urban scene reconstruction aims to reconstruct real-world driving scenes from images and achieve photorealistic rendering [31, 24, 33, 3]. It is crucial for realizing real-world closed-loop evaluation in end-to-end autonomous driving simulators [31, 29]. One of the most popular methods for driving scene reconstruction is the use of Neural Radiance Field (NeRF) [19], an implicit radiance field representation. Although NeRF-based methods [31, 24, 27, 6, 21] have achieved photorealistic rendering, they suffer from slow rendering speed and expensive training cost. Recently, some researchers resort to 3D Gaussian Splatting (3DGS) [9], an explicit radiance field representation, for urban scene reconstruction. 3DGS-based methods [34, 33, 3, 30] use explicit 3D Gaussian primitives to represent 3D scenes and employ a differentiable tile rasterizer for rendering. This enhances editability of driving scenes and achieves real-time photorealistic rendering.
@@
However, existing driving scene reconstruction methods mainly focus on rendering from pinhole cameras and largely neglect fisheye cameras. Compared with pinhole cameras, fisheye cameras provide a wider field of view (FOV) essential for navigation and perception tasks in autonomous driving, especially for near-field sensing and automatic parking [22, 13, 12]. Nevertheless, it is non-trivial to adapt 3DGS to fisheye cameras. For example, the direct application of 3DGS to fisheye cameras is impeded by distortions induced by camera optics, which disrupts the affine transformations of 3D Gaussians. Intuitively, an engineering strategy to tackle this issue is directly rectifying fisheye images and treating them as pinhole images for model training. However, this strategy yields suboptimal rendering quality, particularly in the regions with large field of view angle. Another strategy is rendering multiple cubic images and subsequently amalgamating them to construct fisheye images [11, 25]. Nevertheless, this strategy introduces noticeable artifacts indicative of the stitching process. In fact, how to effectively render driving scenes from fisheye images for simulation remains an open question.
@@
To address the aforementioned challenges, this work presents UniGaussian, a novel 3DGS-based approach that learns a unified Gaussian representation from multiple camera models for urban scene reconstruction in autonomous driving. First, we aim to address the compatibility issue of 3DGS with fisheye cameras. To this end, we propose a novel differentiable rendering method for 3DGS tailored to fisheye cameras. In our method, the distortion of light rays caused by fisheye cameras is translated into the deformation of the radiance field. This is realized through a series of affine transformations, incluiding translation, rotation and stretching, applied to 3D Gaussians. These operations maintain the parallelization strategy and real-time performance of 3DGS, while ensuring differentiability. Second, with the proposed differentiable rendering method, we design a new framework that jointly optimizes a unified 3D Gaussian representation from multiple camera models for driving scene reconstruction. We apply affine transformations on 3D Gaussians to adapt them for different camera models and regularize the shared Gaussians with supervision from different modalities. In this way, our framework learns a unified 3D Gaussian representation with input data from multiple sources and achieves holistic driving scene understanding. As illustrated in Fig. 1, our approach models multiple sensors, including pinhole cameras and fisheye cameras, and multiple modalities, including depth, semantic, normal and optional LiDAR point clouds. Besides, since we use a unified rasterizer for 3D Gaussian rendering, our framework shows adaptability to various differentiable camera models and maintains real-time performance. The framework of the proposed approach is depicted in Fig. 2.
@@
  • We propose a novel differentiable rendering method for 3DGS tailored to driving scene reconstruction from fisheye cameras.
  • We propose a new framework to learn a unified 3D Gaussian representation from multiple camera models for driving scene reconstruction. Our approach achieves holistic driving scene understanding by modeling multiple sensors and modalities.
@@ 2 Related Work
3D Gaussian Splatting.
@@
3DGS [9] is a real-time radiance field rendering approach. It represents a 3D scene with explicit 3D Gaussian primitives. With a differentiable tile-based rasterizer, 3DGS achieves highly effective training and photorealistic real-time rendering. We refer the readers to [9] for more details of 3DGS. With the incredible success achieved by 3DGS, many 3DGS-based methods have emerged with the aim for solving the problems of anti-aliasing [32], acceleration [4], relighting [5], sparse-view synthesis [28], 4D reconstruction [26], etc. Most of these works focus on reconstruction with pinhole camera models only. Recently, Liao et al. [15] adapt 3DGS to fisheye cameras by recalculating 3DGS projection and gradients, but as they point out, their method is only based on ideal camera models with equidistant projection and not for generic fisheye camera models and real large-FOV cameras. This hinders their use in driving scene reconstruction because fisheye cameras in driving scene are usually generic models and have large FOVs. In contrast, our work proposes a new differentiable rendering method for 3DGS tailored to driving scene reconstruction with fisheye cameras. Through a series of affine transformations, our method does not significantly increase GPU memory consumption and maintains real-time rendering. Furthermore, our method has adaptability to various differentiable camera models and can deal with driving scene reconstruction with large FOVs.
Driving Scene Reconstruction.
@@
Typically, autonomous driving scene simulation involves driving scene reconstruction [31, 27, 24] and road asset reconstruction [16, 17]. Our work mainly focuses on driving scene reconstruction from images. Contemporary driving scene reconstruction methods can be mainly categorized into two types, namely NeRF-based methods and 3DGS-based methods. Generally, NeRF-based methods [24, 27, 6, 21, 31] perform dense ray sampling and leverage a multi-layer perception with many fully-connected layers to represent 3D scenes. Although they have shown promising results for photorealistic rendering, they are usually slow for both training and rendering. This hinders their application to real-world autonomous driving simulators. On the other hand, 3DGS-based methods [34, 33, 3, 30] appear to be more efficient for driving scene reconstruction thanks to the efficiency of explicit 3D Gaussian representations and differentiable rasterization. They are able to achieve photorealistic reconstruction and real-time rendering. Nevertheless, existing NeRF-based and 3DGS-based methods neglect the importance of fisheye camera simulation for autonomous driving. In autonomous driving, fisheye cameras are useful for providing a wider field of view, especially for near-field sensing and automatic parking. Our work differs from existing works in that the proposed method is built on a novel differentiable rendering method tailored to fisheye cameras and the proposed framework learns unified 3D Gaussians by modeling multiple sensors and modalities. In contrast, no existing method simulates multiple camera models in a unified framework for driving scene reconstruction.
@@ 3 Methodology
This work aims at realizing real-time photorealistic rendering for driving scene reconstruction from multiple camera models, including pinhole and fisheye cameras. Before delving into driving scene reconstruction from multiple camera models, it is necessary to discuss differentiable rendering of 3DGS for fisheye cameras. Thus, in Sec. 3.1, we briefly introduce mathematical models of fisheye cameras; then, in Sec. 3.2, we present the proposed differentiable rendering of 3DGS for fisheye cameras; finally, in Sec. 3.3, we introduce the proposed UniGaussian framework for driving scene reconstruction.
@@ 3.1 Mathematical Models of Fisheye Cameras
The Kannala-Brandt [8] and MEI [18] mathematical models are two commonly used models for fisheye cameras. Each model consists of two main steps. The first step can be referred to as the “mirror transformation” that characterizes the twisting of light rays induced by the lens/mirrors or the transformation that translates 3D points to new positions by rotating the camera-point vectors towards the optical axis of the camera. The second step is projecting 3D points from new positions onto the image plane. This step utilizes either the camera intrinsic matrix or a pseudo intrinsic matrix.
Let $\theta$ and $\theta_{d}$ be the included angles between camera-point vectors and the optical axis before and after applying the rotation, respectively. Then, the first step is defined as:
@@
$\small\begin{split}\theta_{d}=\arctan r_{d}=\arctan\left(\theta\left(1+k_{1}\theta^{2}+k_{2}\theta^{4}+k_{3}\theta^{6}+k_{4}\theta^{8}\right)\right),\end{split}$ (2)
@@
where $\chi=\frac{\sin{\theta}}{\cos{\theta}+\xi}$ and the twist of light rays is controlled by three parameters $k_{1}$, $k_{2}$ and $\xi$. Next, these translated 3D points are projected onto the image plane with intrinsic parameters $f_{x}$, $f_{y}$, $u_{0}$ and $v_{0}$. In specific, the pseudo intrinsic matrix defined in Eq. (4) is used to project the translated 3D points onto the image plane. Since the focal length $f_{x}$, $f_{y}$ and the gain $\eta$ cannot be estimated independently, $\gamma_{1}$ and $\gamma_{2}$ are used as the pseudo focal length of the MEI model.
$\small\begin{bmatrix}\gamma_{1}&0&u_{0}\\ 0&\gamma_{2}&v_{0}\\ 0&0&1\\ \end{bmatrix}=\begin{bmatrix}f_{x}\eta&0&u_{0}\\ 0&f_{y}\eta&v_{0}\\ 0&0&1\\ \end{bmatrix}.$ (4)
@@ Discussion.
The architecture of the proposed differentiable rendering method is shown in Fig. 3(a). Suppose a driving scene is represented by a set of 3D Gaussians. When rendering from a specific viewpoint, we firstly select all visible 3D Gaussians and compute their colors utilizing spherical harmonics in the world space. Subsequently, we adjust the positions of the selected 3D Gaussians via rotating the camera-Gaussian center vectors towards the optical axis of the camera. Also, the poses of the selected 3D Gaussian are rotated with the camera-Gaussian center vectors. Then, the adjusted 3D Gaussians are compressed in both polar and tangential directions. The scaling factor and the quaternion of the 3D Gaussians are updated via eigendecomposition. Finally, the updated 3D Gaussians are projected onto the image plane to generate the final rendering. Overall, our method is generic to various differentiable camera models and can be used for driving scene reconstruction with large FOVs. Besides, our method does not require large amounts of additional GPU memory so is suitable for processing millions of 3D Gaussians in driving scene reconstruction.
@@
Let $\bm{r}_{g}$ and $\bm{r}_{c}$ be the position of 3D Gaussians and cameras in the world coordinate space respectively and $\bm{r}_{a}$ be the optical axis of cameras. As shown in Fig. 3(b), when performing transformation, the camera-Gaussian center vector needs to be rotated by $\theta_{\Delta}$ around the vector $\bm{r}_{rot}$. Here, $\theta_{\Delta}=\theta_{d}-\theta$ and $\bm{r}_{rot}$ is denoted by:
@@
Corresponding to the compression of FOV, 3D Gaussians needs to be compressed in the polar and tangential directions; otherwise, the adjacent 3D Gaussians will have greater degree of overlap on the splatting plane. To compress 3D Gaussians in the given direction, we use a stretching matrix $\bm{S}$ which is defined as:
$\small\bm{S}(\hat{\bm{n}},k)=\begin{bmatrix}1+(k-1)n_{x}^{2},&(k-1)n_{x}n_{y},&(k-1)n_{x}n_{z}\\ (k-1)n_{x}n_{y},&1+(k-1)n_{y}^{2},&(k-1)n_{y}n_{z}\\ (k-1)n_{x}n_{z},&(k-1)n_{y}n_{z},&1+(k-1)n_{z}^{2}\end{bmatrix},$ (8)
where $\hat{\bm{n}}$ is the stretching direction and $k$ is the stretching ratio. In Fig. 3(c), the blue and green ellipsoids denote 3D Gaussians before and after translation, rotation and stretching. These two ellipsoids have the same tangential field angle $\Delta\phi$. The tangential stretching ratio is computed by $\sin\theta_{d}/\sin\theta$ and the polar stretching ratio is defined as:
@@
$\small\mathcal{U}=((6k_{1}\chi+20k_{2}\chi^{3})(\frac{d\chi}{d\theta})^{2}+(1{+}3k_{1}\chi^{2}{+}5k_{2}\chi^{4})\frac{d^{2}\chi}{d\theta^{2}})(1{+}r_{d}^{2}),$ (16)
@@
Generally, random variables conforming to a Gaussian distribution can retain this property even after applying affine transformations. Since the stretching operation is an affine transformation in our method, the covariance matrix of the resultant Gaussian distribution is defined as:

### v2

@@ Abstract
Urban scene reconstruction is crucial for real-world autonomous driving simulators. Although existing methods have achieved photorealistic reconstruction, they mostly focus on pinhole cameras and neglect fisheye cameras. In fact, how to effectively simulate fisheye cameras in driving scene remains an unsolved problem. In this work, we propose UniGaussian, a novel approach that learns a unified 3D Gaussian representation from multiple camera models for urban scene reconstruction in autonomous driving. Our contributions are two-fold. First, we propose a new differentiable rendering method that distorts 3D Gaussians using a series of affine transformations tailored to fisheye camera models. This addresses the compatibility issue of 3D Gaussian splatting with fisheye cameras, which is hindered by light ray distortion caused by lenses or mirrors. Besides, our method maintains real-time rendering while ensuring differentiability. Second, built on the differentiable rendering method, we design a new framework that learns a unified Gaussian representation from multiple camera models. By applying affine transformations to adapt different camera models and regularizing the shared Gaussians with supervision from different modalities, our framework learns a unified 3D Gaussian representation with input data from multiple sources and achieves holistic driving scene understanding. As a result, our approach models multiple sensors (pinhole and fisheye cameras) and modalities (depth, semantic, norma, and LiDAR point clouds). Our experiments show that our method achieves superior rendering quality and fast rendering speed for driving scene simulation.
Figure 1: An illustration of the proposed approach. UniGaussian reconstructs 3D driving scenes by learning a unified 3D Gaussian representation from multiple input sources. It achieves holistic driving scene understanding and models multiple sensors (pinhole cameras and fisheye cameras) and modalities (semantic, normal, depth, and optional LiDAR point clouds).
[…]
@@ 1 Introduction
Urban scene reconstruction aims to reconstruct real-world driving scenes from images and achieve photorealistic rendering [32, 25, 35, 3]. It is crucial for realizing real-world closed-loop evaluation in end-to-end autonomous driving simulators [32, 30]. One of the most popular methods for driving scene reconstruction is the use of Neural Radiance Field (NeRF) [20], an implicit radiance field representation. Although NeRF-based methods [32, 25, 28, 7, 22] have achieved photorealistic rendering, they suffer from slow rendering speed and expensive training cost. Recently, some researchers have resorted to 3D Gaussian Splatting (3DGS) [10], an explicit radiance field representation, for urban scene reconstruction. 3DGS-based methods [36, 35, 3, 31] use explicit 3D Gaussian primitives to represent 3D scenes and employ a differentiable tile rasterizer for rendering. This enhances editability of driving scenes and achieves real-time photorealistic rendering.
@@
However, existing driving scene reconstruction methods mainly focus on rendering from pinhole cameras and largely neglect fisheye cameras. Compared with pinhole cameras, fisheye cameras provide a wider field of view (FOV) essential for navigation and perception tasks in autonomous driving, especially for near-field sensing and automatic parking [23, 14, 13]. Nevertheless, it is non-trivial to adapt 3DGS to fisheye cameras. For example, the direct application of 3DGS to fisheye cameras is impeded by distortions induced by camera optics, which disrupts the affine transformations of 3D Gaussians. Intuitively, an engineering strategy to tackle this issue is directly rectifying fisheye images and treating them as pinhole images for model training. However, this strategy yields suboptimal rendering quality, particularly in regions with a large field of view angle. Another strategy is rendering multiple cubic images and subsequently amalgamating them to construct fisheye images [12, 26]. Nevertheless, this strategy introduces noticeable artifacts indicative of the stitching process. In fact, how to effectively render driving scenes from fisheye images for simulation remains an open question.
@@
To address the aforementioned challenges, this work presents UniGaussian, a novel 3DGS-based approach that learns a unified Gaussian representation from multiple camera models for urban scene reconstruction in autonomous driving. First, we aim to address the compatibility issue of 3DGS with fisheye cameras. To this end, we propose a novel differentiable rendering method for 3DGS tailored to fisheye cameras. In our method, the distortion of light rays caused by fisheye cameras is translated into the deformation of the radiance field. This is realized through a series of affine transformations, including translation, rotation and stretching, applied to 3D Gaussians. These operations maintain the parallelization strategy and real-time performance of 3DGS, while ensuring differentiability. Second, with the proposed differentiable rendering method, we design a new framework that jointly optimizes a unified 3D Gaussian representation from multiple camera models for driving scene reconstruction. We apply affine transformations on 3D Gaussians to adapt them for different camera models and regularize the shared Gaussians with supervision from different modalities. In this way, our framework learns a unified 3D Gaussian representation with input data from multiple sources and achieves holistic driving scene understanding. As illustrated in Fig. 1, our approach models multiple sensors, including pinhole cameras and fisheye cameras, and multiple modalities, including depth, semantic, normal, and optional LiDAR point clouds. Besides, since we use a unified rasterizer for 3D Gaussian rendering, our framework shows adaptability to various differentiable camera models and maintains real-time performance. The framework of the proposed approach is depicted in Fig. 2.
@@
  • We propose a novel differentiable rendering method for 3DGS tailored to driving scene reconstruction from fisheye cameras. Through a series of affine transformations, our method has adaptability to various fisheye camera models for driving scene reconstruction with large FOVs, does not significantly increase GPU memory consumption and maintains real-time rendering.
  • Built on the proposed differentiable rendering method, we propose to learn unified Gaussian representations from multiple camera models for driving scene reconstruction. Our approach achieves holistic driving scene understanding by modeling multiple sensors and modalities. To the best of our knowledge, no existing driving scene reconstruction method simulates both pinhole and fisheye cameras in a unified framework.
@@ 2 Related Work
2.1 3D Gaussian Splatting
3DGS [10] is a real-time radiance field rendering approach. It represents a 3D scene with a set of explicit 3D Gaussian primitives. Each primitive is defined by a position (mean) $\mu\in\mathbb{R}^{3}$, a covariance matrix $\Sigma\in\mathbb{R}^{3{\times}3}$, an opacity $o\in\mathbb{R}^{1}$ and spherical harmonics (colors). To render images, these 3D Gaussian primitives are projected from 3D space onto 2D image plane to compute the color for each pixel with a differentiable tile-based rasterizer. 3DGS achieves highly effective training and photorealistic real-time rendering. We refer readers to [10] for more details of 3DGS.
@@
With the incredible success achieved by 3DGS, many 3DGS-based methods have emerged with the aim for solving the problems of anti-aliasing [34], acceleration [5], relighting [6], sparse-view synthesis [29], 4D reconstruction [27], etc. Most of these works focus on reconstruction with pinhole camera models only. Recently, Liao et al. [16] adapt 3DGS to fisheye cameras by recalculating 3DGS projection and gradients, but as they point out, their method is only based on ideal camera models with equidistant projection and not for generic fisheye camera models and real large-FOV cameras. This hinders their use in driving scene reconstruction because fisheye cameras in driving scene are usually generic models and have large FOVs. In contrast, our work proposes a new differentiable rendering method for 3DGS tailored to driving scene reconstruction with fisheye cameras. Through a series of affine transformations, our method has adaptability to various fisheye camera models for driving scene reconstruction with large FOVs, does not significantly increase GPU memory consumption and maintains real-time rendering.
2.2 Driving Scene Reconstruction
@@
Typically, autonomous driving scene simulation involves driving scene reconstruction [32, 28, 25] and road asset reconstruction [17, 18]. Our work mainly focuses on driving scene reconstruction from images. Contemporary driving scene reconstruction methods can be mainly categorized into two types, namely NeRF-based methods and 3DGS-based methods. Generally, NeRF-based methods [25, 28, 7, 22, 32] perform dense ray sampling and leverage a multi-layer perception with many fully-connected layers to represent 3D scenes. Although they have shown promising results for photorealistic rendering, they are usually slow for both training and rendering. This hinders their application to real-world autonomous driving simulators. On the other hand, 3DGS-based methods [36, 35, 3, 31] appear to be more efficient for driving scene reconstruction due to the efficiency of explicit 3D Gaussian representations and differentiable rasterization. They are able to achieve photorealistic reconstruction and real-time rendering. Nevertheless, existing NeRF-based and 3DGS-based methods neglect the importance of fisheye camera simulation for autonomous driving. In autonomous driving, fisheye cameras are useful for providing a wider field of view, especially for near-field sensing and automatic parking. Our work differs from existing works in that the proposed method is built on a novel differentiable rendering method tailored to fisheye cameras and the proposed framework learns unified 3D Gaussians by modeling multiple sensors and modalities.
Although we use composite Gaussians as driving scene representation following [35, 36, 31], our approach significantly diffes from [35, 36, 31]. Our approach presents a novel differentiable rendering method tailored to fisheye cameras and constructs a unified framework to learns unified 3D Gaussians by modeling multiple sensors and modalities. In contrast, [35, 36, 31] are designed for modeling driving scene with pinhole cameras only. Specifically, [31] constructs composite street Gaussians as driving scene representation and employs tracking pose optimization to improve foreground object modeling; [35] employs composite Gaussians for holistic scene understanding with a unicycle model for foreground object modeling; [36] presents incremental 3D Gaussians and dynamic Gaussian graph for driving scene reconstruction. None of them simulates multiple camera models in a unified framework for driving scene reconstruction. Note that, post-processing or redistrotion is not a practical solution to 3DGS fisheye rendering in driving scene reconstruction, which yields suboptimal rendering quality in regions with a large FOV angle (see Sec. 4.2 for comparison with 3DGS+Undistort) and prevents each camera model from learning complementary information.
@@ 3 Methodology
This work focuses on driving scene reconstruction from multiple camera models, including pinhole and fisheye cameras. Before delving into driving scene reconstruction from multiple camera models, it is necessary to discuss differentiable rendering of 3DGS for fisheye cameras. Thus, in Sec. 3.1, we briefly introduce mathematical models of fisheye cameras; then, in Sec. 3.2, we present the proposed differentiable rendering of 3DGS for fisheye cameras; finally, in Sec. 3.3, we introduce the proposed UniGaussian framework for driving scene reconstruction.
@@ 3.1 Mathematical Models of Fisheye Cameras
The Kannala-Brandt [9] and MEI [19] mathematical models are two models commonly used for fisheye cameras. Each model consists of two main steps. The first step can be referred to as the “mirror transformation” that characterizes the twisting of light rays induced by the lens/mirrors or the transformation that translates 3D points to new positions by rotating the camera-point vectors towards the optical axis of the camera. The second step is projecting 3D points from new positions onto the image plane. This step utilizes either the camera intrinsic matrix or a pseudo intrinsic matrix.
Let $\theta$ and $\theta_{d}$ be the included angles between the camera-point vectors and the optical axis before and after applying the rotation, respectively. Then, the first step is defined as:
@@
$\begin{split}\theta_{d}=\arctan r_{d}=\arctan\left(\theta\left(1+k_{1}\theta^{2}+k_{2}\theta^{4}+k_{3}\theta^{6}+k_{4}\theta^{8}\right)\right),\end{split}$ (2)
@@
where $\chi=\frac{\sin{\theta}}{\cos{\theta}+\xi}$ and the twist of light rays is controlled by three parameters $k_{1}$, $k_{2}$ and $\xi$. Next, these translated 3D points are projected onto the image plane with intrinsic parameters $f_{x}$, $f_{y}$, $u_{0}$ and $v_{0}$. Specifically, the pseudo intrinsic matrix defined in Eq. (4) is used to project the translated 3D points onto the image plane. Since the focal length $f_{x}$, $f_{y}$ and the gain $\eta$ cannot be estimated independently, $\gamma_{1}$ and $\gamma_{2}$ are used as the pseudo focal length of the MEI model.
$\begin{bmatrix}\gamma_{1}&0&u_{0}\\ 0&\gamma_{2}&v_{0}\\ 0&0&1\\ \end{bmatrix}=\begin{bmatrix}f_{x}\eta&0&u_{0}\\ 0&f_{y}\eta&v_{0}\\ 0&0&1\\ \end{bmatrix}.$ (4)
@@ Discussion.
The architecture of the proposed differentiable rendering method is shown in Fig. 3(a). Suppose a driving scene is represented by a set of 3D Gaussians. When rendering from a specific viewpoint, we first select all visible 3D Gaussians and compute their colors utilizing spherical harmonics in the world space. Subsequently, we adjust the positions of the selected 3D Gaussians via rotating the camera-Gaussian center vectors towards the optical axis of the camera. Also, the poses of the selected 3D Gaussian are rotated with the camera-Gaussian center vectors. Then, the adjusted 3D Gaussians are compressed in both polar and tangential directions. The scaling factor and the quaternion of the 3D Gaussians are updated via eigendecomposition. Finally, the updated 3D Gaussians are projected onto the image plane to generate the final rendering. Overall, our method is generic to various differentiable camera models and can be used for driving scene reconstruction with large FOVs. Besides, our method does not require large amounts of additional GPU memory, so is suitable for processing millions of 3D Gaussians in driving scene reconstruction.
@@
Let $\bm{r}_{g}$ and $\bm{r}_{c}$ be positions of 3D Gaussians and cameras in the world coordinate space, respectively, and let $\bm{r}_{a}$ be the optical axis of cameras. As shown in Fig. 3(b), when performing the transformation, the camera-Gaussian center vector needs to be rotated by $\theta_{\Delta}$ around the vector $\bm{r}_{rot}$. Here, $\theta_{\Delta}=\theta_{d}-\theta$ and $\bm{r}_{rot}$ is denoted by:
@@
Corresponding to the compression of FOV, 3D Gaussians needs to be compressed in the polar and tangential directions; otherwise, adjacent 3D Gaussians will have greater degree of overlap on the splatting plane. To compress 3D Gaussians in the given direction, we use a stretching matrix $\bm{S}$ which is defined as:
$\bm{S}(\hat{\bm{n}},k)=\begin{bmatrix}1+(k-1)n_{x}^{2},&(k-1)n_{x}n_{y},&(k-1)n_{x}n_{z}\\ (k-1)n_{x}n_{y},&1+(k-1)n_{y}^{2},&(k-1)n_{y}n_{z}\\ (k-1)n_{x}n_{z},&(k-1)n_{y}n_{z},&1+(k-1)n_{z}^{2}\end{bmatrix},$ (8)
where $\hat{\bm{n}}$ is the stretching direction and $k$ is the stretching ratio. This stretching matrix $\bm{S}$ is used to scale targets along an arbitrary axis [4]. In Fig. 3(c), the blue and green ellipsoids denote 3D Gaussians before and after translation, rotation, and stretching. These two ellipsoids have the same tangential field angle $\Delta\phi$. The tangential stretching ratio is computed by $\sin\theta_{d}/\sin\theta$ and the polar stretching ratio is defined as:
@@
$\mathcal{U}=((6k_{1}\chi+20k_{2}\chi^{3})(\frac{d\chi}{d\theta})^{2}+(1{+}3k_{1}\chi^{2}{+}5k_{2}\chi^{4})\frac{d^{2}\chi}{d\theta^{2}})(1{+}r_{d}^{2}),$ (16)
@@
Generally, random variables that conform to a Gaussian distribution can retain this property even after applying affine transformations. Since the stretching operation is an affine transformation in our method, the covariance matrix of the resultant Gaussian distribution is defined as: